A U - tube manometer containing mecury gives a height reading of 250mm of mecury when connected to a gas cylinder. If the barometer reading at the same time is 756mm of mecury.What is the absolute pressure of the gas in cylinder?
1
Expert's answer
2019-04-23T11:10:24-0400

The absolute pressure is a sum of atmospheric pressure and gauge pressure. So

"p=p_a+p_g=250+756=1006 \:\rm{mm\; of\; mercury}"

"1006 \:\rm{mm\; of\; mercury}=13600\:\rm{kg/m^3}\times 9.8\:\rm{N/kg}\times 1.006\:\rm{m}""=134000\:\rm{Pa}=134\:\rm{kPa}"